OpenStudy (anonymous):

The legs of a right triangle are 5 cm and 12 cm long. Find the lengths, to the nearest tenth, of the segments into which the bisector of the right angle divides the hypotenuse.

OpenStudy (tkhunny):

It is hoped that you first recognize the Pythagorean Triple, 5-12-13. Check it out. The hypotenuse is 13. Go ahead and draw a picture. Add to your usual right triangle, the andgle bisector of the Right angle and extend the bisector until it intersects the hypotenuse.

OpenStudy (tkhunny):

Label the sides, 5, 12, 13. Don't worry too much about the scale.

OpenStudy (tkhunny):

Where's the 13?

OpenStudy (anonymous):

is it

OpenStudy (anonymous):



OpenStudy (tkhunny):

Yes. Add the little angle to signify a Right angle and then label the angle to the right \(\alpha\). Finally, lable the top angle \(\beta\)

OpenStudy (anonymous):

okay wait so add 12 and 5 together?

OpenStudy (anonymous):

thats why i was confused

OpenStudy (tkhunny):

Why on EARTH would you do that? Just label the angles.

OpenStudy (tkhunny):

Well, do the labeling. Maybe that's why you're struggling. Let the notation keep track of things for you.

OpenStudy (anonymous):

okay I labeled my drawing then what? Im sorry! Im really confused

OpenStudy (tkhunny):

Have a little faith. Is thast lower-right angle labled \(\alpha\) ?

OpenStudy (anonymous):

no?

OpenStudy (tkhunny):

Why not? Label it. Label it "Fred" if you like. Just label it and show me what it is. Label the one at the top, too.

OpenStudy (anonymous):

label them what? thats what I dont get! I wrote the number out so whats the labels for?

OpenStudy (anonymous):

sorry I just really dont know what im doing (obviously)

OpenStudy (tkhunny):



OpenStudy (anonymous):

oh my gosh okay gotcha

OpenStudy (tkhunny):

We just need a way to talk about the parts of the triangle.

OpenStudy (anonymous):

okay gotcha now sorry!

OpenStudy (tkhunny):

What do you know about angle "a"? Homefully, \(\tan(a) = 12/5\). Agreed?

OpenStudy (tkhunny):

* hopefully

OpenStudy (anonymous):

okay

OpenStudy (tkhunny):

Can you find angle 'a'? Tangent button on your calculator, maybe. Whoops! I have to go. Find 'a'. Then angle b = 90 - a. You tell me why. Once you get the measures of angles a and b, we can move on with the angle bisector. Someone will help you, or I'll be back in a while.

OpenStudy (anonymous):

okay thanks....

OpenStudy (tkhunny):

Did you find the measures of angles a and b?

OpenStudy (anonymous):

I dont know haha I finished it already not sure what I even got but I think I got it...